# HyFormer: Revisiting the Roles of Sequence Modeling and Feature Interaction in CTR Prediction

### v1

@@ 4. Experiments
Sequence Modeling Feature Interaction $\uparrow AUC$ Imp. Params ($\times 10^{6}$) FLOPs ($\times 10^{12}$)
@@
LONGER(Chai et al., 2025) Full Transformer(Vaswani et al., 2017) 0.6472 -0.41% 416 6.2
LONGER(Chai et al., 2025) Wukong(Zhang et al., 2024) 0.6465 -0.88% 385 5.2
Full Transformer(Vaswani et al., 2017) RankMixer(Tolstikhin et al., 2021; Zhu et al., 2025) 0.6481 +0.20% 388 6.6
Full Transformer(Vaswani et al., 2017) Full Transformer(Vaswani et al., 2017) 0.6474 -0.27% 418 9.3
Full Transformer(Vaswani et al., 2017) Wukong(Zhang et al., 2024) 0.6468 -0.68% 387 8.3
@@
MTGR/OneTrans (w/ LONGER)(Han et al., 2025; Zhang et al., 2025) 0.6480 +0.14% 406 6.6
MTGR/OneTrans (w/ Full Transformer)(Han et al., 2025; Zhang et al., 2025) 0.6483 +0.34% 450 21.9
HyFormer (Ours) 0.6489 +0.74% 418 3.9
@@ 4.1.3. Evaluation Metrics
For offline evaluation, we employ Query-level AUC (Area Under the Curve) which calculates the AUC(Hand and Till, 2001) for samples within each query and then averages the results across all queries. To assess the relative efficacy, we adopt the relative improvement (Imp.)(Yan et al., 2014) as our key metric for measuring gains in AUC. We also report the number of dense parameters and training FLOPs, with the latter computed using a batch size of 2048.
@@ 4.2. Overall Performance
Table 2 presents an ablation study on the primary contributors to HyFormer’s performance improvement. First, we ablate the components of the query. The HyFormer query is generated from three sources: global non-sequential features, multiple sequences pooling tokens, and the original target features. Experiments show that reverting the query to its original, target-feature-only state severely limits subsequent deep feature interaction, causing a 0.34% AUC decline. Removing the cross-sequence pooling tokens from the full query also leads to a 0.20% AUC loss, confirming that inter-sequence interaction contributes meaningfully within the HyFormer structure.
Second, we evaluate the overall architectural change. Restoring the baseline architecture (LONGER + RankMixer), which applies sequential modeling followed by separate feature interaction, shows that even with enriched query information, the lack of deepened interaction caps the gains, yielding only a 0.14% AUC improvement(-0.74% vs -0.60%). In contrast, within the HyFormer framework which is designed to strengthen interaction throughout the model, expanding query information delivers a significantly larger 0.34% AUC gain.
@@
Configuration $\uparrow AUC$ Imp. Params ($\times 10^{6}$) FLOPs ($\times 10^{12}$)
@@
Query w/o Seq Pooling Tokens 0.6486 -0.20% 415 3.9
Query w/o Nonseq and Seq Pooling Tokens 0.6484 -0.34% 414 3.8
@@
HyFormer w/o Global Tokens 0.6484 -0.34% 414 3.8
BaseArch w/ Global Tokens 0.6480 -0.60% 505 3.6
BaseArch w/o Global Tokens 0.6478 -0.74% 387 3.5
@@
HyFormer + Merge Seq 0.6485 -0.27% 397 3.9
Third, we conduct an ablation study on the multi-sequence modeling strategy within HyFormer. Two principal paradigms exist for handling multiple sequences: merging sequences into one via dimension alignment and concatenation for joint modeling, or keeping sequences separate and modeling them independently. HyFormer adopts the latter approach, using distinct query tokens for each sequence. In our experiments, it is observed that the sequence merging and query sharing resulted in a significant AUC loss of 0.27%. Consequently, this presents the advantages of the HyFormer in expanding queries and enabling broader feature interaction. Additionally, merging forces disparate sequences to share global tokens, ignoring their distinctiveness. The resulting representations capture far less differentiated information than HyFormer’s separate modeling of each sequence. We speculate that this inherent limitation of sequence merging also partly explains why models like MTGR and OneTrans underperform compared to HyFormer.
@@ 4.3.2. Sparse Dim
224 0.6507 +0.53% 0.26%
We also analyzed how model performance varies with the expansion of the sequence token input dimension (sparse embedding dim), i.e., the richness of sequence side information. Our experiments show that, regardless of sequence length, enriching sequence side information consistently brings greater benefits to the HyFormer framework than to the baseline LONGER + RankMixer framework. As shown in Table 3, for sequences of length 1000, expanding the sparse dimension width from the original 64 dimensions with three side information types (item ID, search query textnet classification, and timestamp) to 224 dimensions with seven types (adding search query ID, author ID, event ID, and playtime) yielded a AUC imp. of 0.41% for the baseline, compared to a 0.54% gain for HyFormer. The improvement for HyFormer is significantly larger, a trend that holds across other sequence lengths in experiments. Furthermore, the performance gap between HyFormer and the BaseArch widens as sequences grow longer, with the additional gain from dimension expansion increasing from 0.13% at 1k sequence length to 0.26% at 3k.

### v2

@@ 4. Experiments
[…]
@@
LONGER(Chai et al., 2025) Full Transformer(Vaswani et al., 2017) 0.6472 -0.09% 416 6.2
LONGER(Chai et al., 2025) Wukong(Zhang et al., 2024) 0.6465 -0.20% 385 5.2
Full Transformer(Vaswani et al., 2017) RankMixer(Tolstikhin et al., 2021; Zhu et al., 2025) 0.6481 +0.05% 388 6.6
Full Transformer(Vaswani et al., 2017) Full Transformer(Vaswani et al., 2017) 0.6474 -0.06% 418 9.3
Full Transformer(Vaswani et al., 2017) Wukong(Zhang et al., 2024) 0.6468 -0.15% 387 8.3
@@
MTGR/OneTrans (w/ LONGER)(Han et al., 2025; Zhang et al., 2025) 0.6480 +0.03% 406 6.6
MTGR/OneTrans (w/ Full Transformer)(Han et al., 2025; Zhang et al., 2025) 0.6483 +0.08% 450 21.9
HyFormer (Ours) 0.6489 +0.17% 418 3.9
@@ 4.1.3. Evaluation Metrics
For offline evaluation, we employ Query-level AUC (Area Under the Curve) which calculates the AUC(Hand and Till, 2001) for samples within each query and then averages the results across all queries. We also report the number of dense parameters and training FLOPs, with the latter computed using a batch size of 2048.
@@ 4.2. Overall Performance
Table 2 presents an ablation study on the primary contributors to HyFormer’s performance improvement. First, we ablate the components of the query. The HyFormer query is generated from three sources: global non-sequential features, multiple sequences pooling tokens, and the original target features. Experiments show that reverting the query to its original, target-feature-only state severely limits subsequent deep feature interaction, causing a 0.08% AUC decline. Removing the cross-sequence pooling tokens from the full query also leads to a 0.05% AUC loss, confirming that inter-sequence interaction contributes meaningfully within the HyFormer structure.
Second, we evaluate the overall architectural change. Restoring the baseline architecture (LONGER + RankMixer), which applies sequential modeling followed by separate feature interaction, shows that even with enriched query information, the lack of deepened interaction caps the gains, yielding only a 0.03% AUC improvement(-0.14% vs -0.17%). In contrast, within the HyFormer framework which is designed to strengthen interaction throughout the model, expanding query information delivers a significantly larger 0.08% AUC gain.
@@
[…]
@@
Query w/o Seq Pooling Tokens 0.6486 -0.05% 415 3.9
Query w/o Nonseq and Seq Pooling Tokens 0.6484 -0.08% 414 3.8
@@
HyFormer w/o Global Tokens 0.6484 -0.08% 414 3.8
BaseArch w/ Global Tokens 0.6480 -0.14% 505 3.6
BaseArch w/o Global Tokens 0.6478 -0.17% 387 3.5
@@
HyFormer + Merge Seq 0.6485 -0.06% 397 3.9
Third, we conduct an ablation study on the multi-sequence modeling strategy within HyFormer. Two principal paradigms exist for handling multiple sequences: merging sequences into one via dimension alignment and concatenation for joint modeling, or keeping sequences separate and modeling them independently. HyFormer adopts the latter approach, using distinct query tokens for each sequence. In our experiments, it is observed that the sequence merging and query sharing resulted in a significant AUC loss of 0.06%. Consequently, this presents the advantages of the HyFormer in expanding queries and enabling broader feature interaction. Additionally, merging forces disparate sequences to share global tokens, ignoring their distinctiveness. The resulting representations capture far less differentiated information than HyFormer’s separate modeling of each sequence. We speculate that this inherent limitation of sequence merging also partly explains why models like MTGR and OneTrans underperform compared to HyFormer.
@@ 4.3.2. Sparse Dim
224 0.6507 +0.12% +0.06%
[…]